OpenStudy (mathsucks321):

Sarah needs to have a plumber come to her house and fix a leaky faucet in her  bathroom. The plumber charges $50 to come to her home and an additional $30.75  every hour he works on the job. Sarah only has $234.50 in her savings for the  repair. What is the maximum amount of hours that Sarah can have the plumber work  on her faucet? A 9 B 7 C 6 D 8

OpenStudy (mathsucks321):

@mathstudent55

OpenStudy (berrymox):

234.50 = 50+30.75h

OpenStudy (mathsucks321):

Explain?

OpenStudy (berrymox):

One time charge of 50, plus 30.75 per every hour is equal to the max money Sarah has to pay with

OpenStudy (mathsucks321):

How would I get h

OpenStudy (berrymox):

h is just a variable for hours, solve for h

OpenStudy (berrymox):

Or do you not know how?

OpenStudy (mathsucks321):

Wait ill try

OpenStudy (mathsucks321):

I got 6 :/

OpenStudy (berrymox):

yup

OpenStudy (mathsucks321):

Ty

OpenStudy (berrymox):

6 is not just 6, but the number of hours

OpenStudy (mathsucks321):

I know :)

OpenStudy (mathstudent55):

Plumber cost: fixed $50 plus $30.75 per hour If the plumber works h hours, the hourly fee will be 30.75h The plumber cost, then, is 30.75h + 50 She has 234.50 The cost expression must be less than or equal to 234.50 \(30.75h + 50 \le 234.50\) Solve for h. Subtract 50 from both sides: \(30.75h \le 184.50\) Divide both sides by 30.75 \(h \le 6\) Answer: she can hire the plumber for at most 6 hours.

OpenStudy (mathsucks321):

Thank you both!!

OpenStudy (mathstudent55):

You're welcome.